Project

General

Profile

11222013

Nightly Observing Plan for MJD 56619 (starting evening of Nov 22, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Arnau Pujol Kevin Reil Marisa March
ext. 511 ext. 513 ext. 516 or 630 338 9267

CTIO staff (Telescope Operators)

Day Support Night Assistant
Manuel Hernandez Alberto Alvarez

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56619, date 20131122, Site CTIO, UT-offset 3 hr
Current local time 17:50 CLST     UT 20:50     SidTim 20:16
Sunset     20:31 CLST  23:31 UT   SidTim  344 deg  
Eve. -10   21:10 CLST  00:10 UT   SidTim  354 deg  Zeropoint; standards
Eve. -14   21:32 CLST  00:32 UT   SidTim  359 deg  Begin obstac
Eve. -18   21:55 CLST  00:55 UT   SidTim    5 deg  Last chance standards
Moonrise   00:20 CLST  03:20 UT   SidTim   42 deg  
Midpoint   01:29 CLST  04:29 UT   SidTim   58 deg  
Morn -18   05:03 CLST  08:03 UT   SidTim  112 deg  Begin astronomical twilight
Morn Std   05:26 CLST  08:26 UT   SidTim  118 deg  End obstac; begin standards
Morn -10   05:48 CLST  08:48 UT   SidTim  123 deg  
Sunrise    06:27 CLST  09:27 UT   SidTim  133 deg  
Moon position (deg): RA 124.0, Dec 15.1 Phase: 0.650

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Clear, some cloud on horizon.

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards.
  3. Nightly out of focus image. script to use: defocus_nightly.json (right after last standard, no break)
  4. ObsTac.
  5. End of night standards.
  6. End of night calibrations. if start of night calibrations good use: postnight_calibration.json, otherwise use postnight_zeros_flats.json

Start of Night Standards MANUAL SLEW TO THESE FIELDS Wind screen down.
UT Date = '2013/11/23 00:10' Local time = 21:10
SDSSJ0320-0000 ra 03:20:00 dec +00:00:00 secz 2.1 script to use: DES_std_sdssj0320_Yzirg
SDSSJ0200-0000 ra 02:00:00 dec +00:00:00 secz 1.4 script to use: DES_std_sdssj0200_Yzirg
SDSSJ0000-0000 ra 00:00:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0000_Yzirg

End of Night Standards MANUAL SLEW TO THESE FIELDS Check Wind screen down.
UT Date = '2013/11/23 08:26' Local time = 05:26
C26202/HST ra 03:32:30 dec -27:46:05 secz 1.8 script to use: DES_std_c26202_grizY
SDSSJ1048-0000 ra 10:48:00 dec +00:00:00 secz 1.6 script to use: DES_std_sdssj1048_grizY
SDSSJ0843-0000 ra 08:43:00 dec +00:00:00 secz 1.2 script to use: DES_std_sdssj0843_grizY

Notes from 4pm Meeting

(Thanks to Tom for minutes)

Run Manager's Meeting 11/22/2013
--------------------------------
Kevin, Marisa, Tom, Eric, Manuel, Liz, Josh
Don, Robert, Margaret, Michael

Camera and Telescope (Manuel)
-----------------------------
Needed to reboot server 1,2,3, readout, pipeline, and system1
All rebooting is done. Monsoon is powered up. Filter-i is in.
Vsub is off.

Test zeros images needed. They show OK. Some test flats coming.

Telemetry Hardware
-------------------
Kevin stays up for the weekend, then La Serena for a night or two,
then back up through early December. He's going to bug Michael
about accelerometers.

3D anemometer will be there in a few weeks. Some 6 RTDs are
curled up available in the

Last Night
----------
4 hrs of good observing but at 04:00 UTC the computers crapped out.
server1 and server3 needed to be rebooted in a certain order.
Liz to see if there is anything in the logs - there was some
telemetry missing about some of the backplanes early on, maybe
starting early in the week.

Observer 2 got bogged-down - mouse slowed, looked like the kinds
of problems that lead to rebooting Observer2 last week.
The O1 console showed exposures not advancing - all the displays
got stuck, except the telemetry continued to cycle.

At that point the observers realized there was a bigger problem.
Of course the alarms were stuck t00. Marisa decided to shutdown and
restart sispi, but she couldn't get a rise out of the machine to do it.
At that stage David R started rebooting O1, but that failed
"configuration". That didn't help. David started looking at the
servers - lots of error messages on server1 and also server3.
But these computers were slow to boot up because of a disk check.

It was decided to safeguard the camera - close dome, shut mirror covers.
Monsoon crates powered down (contact of Alistair). Shutter was
reset, guarenteeing it's closed.
When came back on i-band filter was still in place.

Alistair mentioned server1 and 3 lost connection to KVM switch in the rack.
KVM = keyboard video mouse. Probably a side-effect.
Server1 is the critical macine.
It took an hour to realize server1 had gone down.
There are availalbe tools in the world to monitor the computing systems.
Probably it runs in La Serena. Probably there should be such things in place
visibly on the mountain.

It would help to have hardcopy with info on how to make the camera safe.
There should be a better standard procedure.

But we did get some data.
X1 was retaken. And some WF riz sub-arcsecond.

Eric will nightsummary from last night.

DESDM
-----
Data from last night made it.
Last image from nightsum was 257298.

WF DQ from 11/20 is available from Margaret.
Not yet started processing data from last night - not going
to start before the NCSA power outage. We enter a 4 night
C.O. period.

Tonight
-------
Looks good at CTIO.
SN is in good shape - some fields at 4 nights tonight,
so if seeing is poor we might get some more.

gri when Moon is down
izY when the Moon is up